Two individuals, A and B, are finalists for a chess championship. They will play a sequence of games, each of which can resulNEED LAST 2
Answer #1

a)P(A wins ) =P(wins first 5 games) =0.3^5=0.00243

b)P(ends in 6 games) =P(A wins first 5 games)+P(B wins first 5 games)=0.3^5+0.2^5 =0.00275

c)

Use a random number table, and let the numbers 0-2 represent A winning the game and the numbers 3-4 represent 8 winning, then